OpenStudy (anonymous):

passes through (-1,1) slope -1

OpenStudy (butterflydreamer):

do you know the point-slope formula? :)

OpenStudy (anonymous):

YES

OpenStudy (butterflydreamer):

thennn you would want to use that ! :)

OpenStudy (anonymous):

OKA for this question it is point slope formula and what about this has y intercept 4 and x intercept -1

OpenStudy (anonymous):

what formula for this one

OpenStudy (butterflydreamer):

yaas ^_^. Hmm.. the second one you would use the same thing. But you should remember that y intercept at 4 means (0,4) and x intercept at -1 means (-1,0) :)

OpenStudy (anonymous):

oh same formula for that one too you mean y+=y1+m(x-x1)

OpenStudy (anonymous):

we will plug in the value just?

OpenStudy (anonymous):

and what about this one slope of 1/2 and y intercept -3

OpenStudy (butterflydreamer):

i think so.. You would want it in the form y= mx + b where "b" is your y intercept -3. :)

OpenStudy (anonymous):

oka and passes through 5,-1 parallel to 2x+5y=15

OpenStudy (butterflydreamer):

you'd have to find the gradient of 2x+5y = 15 first. Then use that gradient and your point (5,-1) in the point slope formula again.

OpenStudy (anonymous):

oh oka thank you so so much

OpenStudy (butterflydreamer):

no problem :)!

OpenStudy (anonymous):

i need help with induction too

OpenStudy (anonymous):

;(

OpenStudy (anonymous):

i have calculus final paper on tuesday 1st semester of biotechnology i am in

OpenStudy (butterflydreamer):

if you have questions, you should just post it on here and i'm sure someone will be able to help you out :P Good luck with your finals!

OpenStudy (anonymous):

thankyou so so much ! and i will but m afraid some explanations of questions cant be understand through chatting . i might look into tutorials on youtube

OpenStudy (anonymous):

thankyou so so much ! and i will but m afraid some explanations of questions cant be understand through chatting . i might look into tutorials on youtube